Solve the system of fuctions by graphing then classify the system.
f(x)= x+1
g(x)= -8x+10

Answer by ewatrrr(24785)


Hi,
Graphing
f(x)= x+1 Pts (0,1), (-1,0), (1,2)
g(x)= -8x+10 PTs (0,10), (5/4,0), (1,2)
Lines intersect at Pt(1,2), the solution for this system of equations
Note: when the lines intersect f(x) = g(x)
x+ 1 = -8x + 10
9x = 9
x = 1 Substituting x=1 into either of the equations, will give
the same f(x) and g(x) value = 2, thus the Pt of intersection (1,2)